# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

### v2

@@ A.3 Proof of Theorem 1
$(1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n)^{2}$ $=\left(1-\mathsf{Tr}((\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})/n\right)^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda}{\lambda+D_{ii}^{2}}\right)^{2}=\lambda^{2}v_{\mathbf{D}}(-\lambda)^{2}$
@@
We claim that $\limsup_{n\to\infty}\|\mathbf{T}_{4}\|/\sqrt{n}$ is bounded. This follows from the bound above, the law of large numbers applied to $\boldsymbol{\epsilon}$, and the assumption on $\mathbf{D}$. Hence the bound above is summable in $n$ once more, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. The convergence for $T_{3}$ follows directly from [7, Lemma C.3]
A.4 Proof of Theorem 2
@@
The normalization assumption (Assumption A5) ensures that the prefactor is finite. Without further loss of generality, we can assume it is 1. In fact it is sufficient for all proofs to simply assume that lim $\limsup\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\leq C$, but assuming it is constant simplifies notation.
For intuition why this term is bounded, note that $\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})=\mathsf{Tr}(\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}})$, which is the covariance matrix of $n^{\prime}$ samples normalized by $\sqrt{n}$ (the reason for this is given in Remark 2 – since $\boldsymbol{\beta}$ is scaled up by $\sqrt{n}$ all data must be scaled down by $\sqrt{n}$). Hence we expect this trace to be order $n^{\prime}p/n$, which exactly cancels.
@@
This yields pointwise convergence for the bias. To strengthen this to uniform convergence, we show that the derivative of the difference $\mathbf{b}^{\top}\mathbf{P}\mathbf{b}-\frac{1}{p}\mathsf{Tr}(\mathbf{P})$ is almost surely bounded on $[\lambda_{1},\lambda_{2}]$:
@@
which is indeed almost surely bounded on $[\lambda_{1},\lambda_{2}]$ in the limit by assumption on $\|\mathbf{D}\|$. Similarly, $\frac{\mathrm{d}}{\mathrm{d}\lambda}(\frac{1}{p}\mathsf{Tr}(\mathbf{P}))$ is also bounded on this interval by the same quantity, meaning the difference is Lipschitz. Hence one can discretize the interval $[\lambda_{1},\lambda_{2}]$ into a sufficiently fine grid. Pointwise convergence on every point holds at every point on the grid, then the Lipschitz condition assures that the convergence is uniform.
@@
Then by [7, Lemma C.3], this term satisfies
@@ A.5 Proof of Corollary 3
[…]
@@ A.6 Proof of Lemma 1
A.7 Proof of Corollary 2
A.7.1 Nondegeneracy on the spectrum of $\mathbf{D}$
@@
    There exists $i_{*}\in\{2,\dots,L\}$ such that lim $\liminf|a_{i_{*}}b_{1}-a_{1}b_{i_{*}}|>C$ almost surely for some $C>0$.
@@
As motivation, we first discuss the simpler case when $L=2$, as this has connections to the assumptions required for consistent estimation in [23]. In this setting, the estimator for $\sigma^{2}$ becomes
@@
lim $\liminf\left|a_{1}b_{2}-a_{2}b_{1}\right|=\liminf\left|\frac{1}{p}\sum_{i=1}^{p}\frac{\lambda^{2}D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}-\left(\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)\right|>c>0$ (22)
almost surely for some $c$. This is exactly the second part of our nondegeneracy assumption. Furthermore, this is the exact analog of [23, Assumption 7] in our setting. The above argument thus establishes that consistent estimation is possible if one assumes equation (22).
@@
A.7.2 Proof
Having now overviewed a specific case of our method that yields consistent estimation through the use of an additional assumption already used in the literature, we now turn to the general setting. Note that in the general setting, the condition is weaker – we only require that lim $\liminf|a_{1}b_{i}-a_{i}b_{1}|>C>0$ holds for one of the $i$, meaning one loses nothing by using multiple $\lambda_{i}$, and in fact in practice we observe this results in improved stability.
@@ A.7.2 Proof
As a result, we will require an that the coefficient of both error terms is uniformly bounded, which will yield that our estimators are consistent, meaning it suffices to show
lim $\limsup\max\left(\left|\frac{\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}\right|,\left|\frac{\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}\right|\right)<C$ (23)
@@
lim $\limsup\frac{\left|\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})\right|}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot C^{\prime\prime}\cdot 2\cdot C^{\prime\prime}$
@@
lim $\limsup\frac{\left|\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})\right|}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot 2\cdot C^{\prime\prime},$
@@ A.8 Proof of Theorem 3
Most computations are analogous to the proof of Theorem 2. We first compute $\mathbb{E}[\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\mid\mathbf{X},\mathbf{y}]$ using [31, Lemma 4]. We find that it is equal to $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$. Call this quantity $\mathbf{C}$ for convenience. We then move to computing the risk.
@@
$\mathbb{E}\left[\|\widetilde{\mathbf{X}}(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}})\|_{2}^{2}\right]=(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}})^{\top}\left(\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}\right)(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}}).$ (25)
@@
On the other hand, the second term is order $1$ and hence vanishes when divided by $n$. The convergence is then uniform because the function is monotonic in $\lambda$. It is almost sure through, for instance, using [7, Lemma C.3]. Note that the form of $\mathcal{V}_{\mathbf{X}}$ contains an extra
@@
Recall we are dividing everything by $n$. Hence if we show that $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}\xrightarrow{a.s.}0$, then we will be done with a union bound. We rewrite $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|=\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}$ and can now apply [7, Lemma C.3] to find
@@
Thus $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|^{2}/n$ converges a.s. to 00, and thus by continuous mapping so does $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}$.
A.9 Proof of Lemma 4
The following is taken from [23]. We note that in this work, the errors are assumed Gaussian. We will mark which ones require this and how we believe they can be relaxed. First, the consistency of $\alpha_{i}$ is immediate from the first result of [23, (40)], where we note that the proof does not require Gaussianity of the errors. The one step in the proof where care must be made is in controlling
@@ A.9 Proof of Lemma 4
but we note that this can be done through [7, Lemma C.3].
Then the reduced model is right-rotationally invariant by [23, p. 58, (124)]. Showing that the reduced model is right-rotationally invariant requires Gaussianity and is what leads to the added assumption of Lemma 4. Without the Gaussian assumption, the errors in the reduced model will not be i.i.d. However, in general we only require that certain inner products concentrate. We believe this will still hold in the reduced model when one explicitly tracks the form of the error.
A.10 Proof of Corollary 5
@@
A.11 Weak Convergence of Spectral Measure
Remark 9.
In cases such as the i.i.d. case, the diagonal entries of $\mathbf{D}$ converge weakly almost surely to some limiting spectral distribution. Under the following assumption, one can, instead of stating results in terms of the difference in the predicted and actual risks converging almost surely to zero, as in Corollary 3, instead write them as converging almost surely to the Stieltjes transforms taken over the limiting measure:
  • (A*)
    Almost surely, the diagonal entries of $\mathbf{D}$ converge in the Wasserstein-2 ($W_{2}$) distance to a measure $\mathsf{D}$, where $\mathsf{D}^{2}$ has non-zero mean and compact support $\mathrm{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. See Appendix LABEL:app:w2 for details on Wasserstein-2 convergence.
Lemma 10.
Under the assumptions of Theorem 1 and Assumption (A*), then one has, for any $\lambda_{1},\lambda_{2}$ satisfying $0<\lambda_{1}<\lambda_{2}$ and any two consistent estimators $\hat{r}^{2},\hat{\sigma}^{2}$,
[…]
As usual, if the two estimators are instead strongly consistent, then the above convergence is almost sure.
This is then similar in flavor to [29, Theorem 4.1].
Appendix B Miscellaneous Details
B.1 Experimental details
@@
B.2 Semi-real experimental details
Figure 3: TR (Tuned Risk): test risk attained by the value of $\lambda$ chosen through the given CV method. ER (Estimated Risk): test risk estimated by the CV method. MR (MinRisk): attainable value of out-of-sample risk. Plot format same as in main manuscript. Each plot shows 10 resamplings of noise. Dashed blue line shows average actual MSE (over resamplings). Black line shows average CV risk curve for that CV method. Standard deviation of all quantities shown in parentheses.
@@ B.2 Semi-real experimental details
  5. 5.
    While in the other plots in the main manuscript, we only resampled $\epsilon$, the choice of $\beta$ affects the loss curves more here. We resample $\beta$ four times (referring to Trials 1 through 4 in Figure 3), where we resample $\epsilon$ ten times per trial. In the main manuscript, only one trial was shown.
B.3 Proof of right-rotational invariance
@@
Matrix normals: Recall that $\boldsymbol{\Sigma}^{\mathrm{col}}$ is drawn from an inverse Wishart distribution with identity scale and $(1+\delta)p$ degrees of freedom; this can be generated by $(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$, where $\mathbf{G}_{2}\in\mathbb{R}^{(1+\delta)p\times p}$ has i.i.d. standard Gaussian entries (since the inverse Wishart has identity scale) [11, Appendix A]. A matrix normal with row covariance $\boldsymbol{\Sigma}^{\text{row}}$ and column covariance $\boldsymbol{\Sigma}^{\text{col}}=(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$ can then be generated by $(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}$ (with $\mathbf{G}$ again having i.i.d. Gaussian entries and being independent of $\mathbf{G}_{2}$). The claim then follows from the following. For any fixed $\mathbf{O}\in\mathbb{O}_{p}$,
$(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}\mathbf{O}$ =$\stackrel{{\scriptstyle d}}{{=}}(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}\mathbf{O}(\mathbf{O}\mathbf{G}_{2}^{\top}\mathbf{G}_{2}\mathbf{O}^{\top})^{-1/2}\mathbf{O}$

### v3

@@ 7.3 Proof of Theorem 1
$(1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n)^{2}$ $=\left(1-\mathsf{Tr}((\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})/n\right)^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda}{\lambda+D_{ii}^{2}}\right)^{2}=\gamma^{2}\lambda^{2}v_{\mathbf{D}}(-\lambda)^{2}$
@@
We claim that $\limsup_{n\to\infty}\|\mathbf{T}_{4}\|/\sqrt{n}$ is bounded. This follows from the bound above, the law of large numbers applied to $\boldsymbol{\epsilon}$, and the assumption on $\mathbf{D}$. Hence the bound above is summable in $n$ once more, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. The convergence for $T_{3}$ follows directly from (Dobriban and Wager, , 2018, Lemma C.3)
7.4 Proof of Theorem 2
@@
The normalization assumption (Assumption A5) ensures that the prefactor is finite. Without further loss of generality, we can assume it is 1. In fact it is sufficient for all proofs to simply assume that $\limsup\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\leq C$, but assuming it is constant simplifies notation.
[…]
@@
This yields pointwise convergence for the bias. To strengthen this to uniform convergence, we show that the derivative of the difference $\mathbf{b}^{\top}\mathbf{P}\mathbf{b}-\frac{1}{p}\mathsf{Tr}(\mathbf{P})$ is almost surely bounded on $\lambda_{1},\lambda_{2}]$:
@@
which is indeed almost surely bounded on $\lambda_{1},\lambda_{2}]$ in the limit by assumption on $\|\mathbf{D}\|$. Similarly, $\frac{\mathrm{d}}{\mathrm{d}\lambda}(\frac{1}{p}\mathsf{Tr}(\mathbf{P}))$ is also bounded on this interval by the same quantity, meaning the difference is Lipschitz. Hence one can discretize the interval $\lambda_{1},\lambda_{2}]$ into a sufficiently fine grid. Pointwise convergence on every point holds at every point on the grid, then the Lipschitz condition assures that the convergence is uniform.
@@
Then by (Dobriban and Wager, , 2018, Lemma C.3), this term satisfies
@@ 7.5 Proof of Corollary 3
$\sup_{\lambda>\lambda_{0}}\left|(\hat{r}^{2}-r^{2})\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)+(\hat{\sigma}^{2}-\sigma^{2})(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))\right|\xrightarrow{p}0.$ (21)
@@ 7.6 Proof of Lemma 1
7.7 Proof of Corollary 2
7.7.1 Nondegeneracy on the spectrum of $\mathbf{D}$
@@
    There exists $i_{*}\in\{2,\dots,L\}$ such that $\liminf|a_{i_{*}}b_{1}-a_{1}b_{i_{*}}|>C$ almost surely for some $C>0$.
@@
As motivation, we first discuss the simpler case when $L=2$, as this has connections to the assumptions required for consistent estimation in Li and Sur, (2023). In this setting, the estimator for $\sigma^{2}$ becomes
@@
$\liminf\left|a_{1}b_{2}-a_{2}b_{1}\right|=\liminf\left|\frac{1}{p}\sum_{i=1}^{p}\frac{\lambda^{2}D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}-\left(\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)\right|>c>0$ (22)
almost surely for some $c$. This is exactly the second part of our nondegeneracy assumption. Furthermore, this is the exact analog of (Li and Sur, , 2023, Assumption 7) in our setting. The above argument thus establishes that consistent estimation is possible if one assumes equation (22).
@@
7.7.2 Proof
Having now overviewed a specific case of our method that yields consistent estimation through the use of an additional assumption already used in the literature, we now turn to the general setting. Note that in the general setting, the condition is weaker – we only require that $\liminf|a_{1}b_{i}-a_{i}b_{1}|>C>0$ holds for one of the $i$, meaning one loses nothing by using multiple $\lambda_{i}$, and in fact in practice we observe this results in improved stability.
@@ 7.7.2 Proof
As a result, we will require that the coefficient of both error terms is uniformly bounded, which will yield that our estimators are consistent, meaning it suffices to show
$\limsup\max\left(\left|\frac{\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}\right|,\left|\frac{\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}\right|\right)<C$ (23)
@@
$\limsup\frac{\left|\sum_{i=1}^{L}(a_{i}^{-1}-a_{1}^{-1})(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})\right|}{\sum_{i=1}^{L}(b_{i}a_{i}^{-1}-b_{1}a_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot C^{\prime\prime}\cdot 2\cdot C^{\prime\prime}$
@@
$\limsup\frac{\left|\sum_{i=1}^{L}(b_{i}^{-1}-b_{1}^{-1})(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})\right|}{\sum_{i=1}^{L}(a_{i}b_{i}^{-1}-a_{1}b_{1}^{-1})^{2}}$ $\stackrel{{\scriptstyle a.s.}}{{\leq}}\frac{1}{C^{\prime}}L\cdot 2\cdot C^{\prime\prime},$
@@ 7.8 Proof of Theorem 3
Most computations are analogous to the proof of Theorem 2. We first compute $\mathbb{E}[\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\mid\mathbf{X},\mathbf{y}]$ using (Rangan et al., , 2019, Lemma 4). We find that it is equal to $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$. Call this quantity $\mathbf{C}$ for convenience. We then move to computing the risk.
@@
[…]
@@
On the other hand, the second term is order $1$ and hence vanishes when divided by $n$. The convergence is then uniform because the function is monotonic in $\lambda$. It is almost sure through, for instance, using (Dobriban and Wager, , 2018, Lemma C.3). Note that the form of $\mathcal{V}_{\mathbf{X}}$ contains an extra
@@
Recall we are dividing everything by $n$. Hence if we show that $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}\xrightarrow{a.s.}0$, then we will be done with a union bound. We rewrite $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|=\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}$ and can now apply (Dobriban and Wager, , 2018, Lemma C.3) to find
@@
Thus $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|^{2}/n$ converges a.s. to 0, and thus by continuous mapping so does $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}$.
7.9 Proof of Lemma 4
The following is taken from Li and Sur, (2023). We note that in this work, the errors are assumed Gaussian. We will mark which ones require this and how we believe they can be relaxed. First, the consistency of $\alpha_{i}$ is immediate from the first result of (Li and Sur, , 2023, (40)), where we note that the proof does not require Gaussianity of the errors. The one step in the proof where care must be made is in controlling
@@ 7.9 Proof of Lemma 4
but we note that this can be done through (Dobriban and Wager, , 2018, Lemma C.3).
Then the reduced model is right-rotationally invariant by (Li and Sur, , 2023, p. 58, (124)). Showing that the reduced model is right-rotationally invariant requires Gaussianity and is what leads to the added assumption of Lemma 4. Without the Gaussian assumption, the errors in the reduced model will not be i.i.d. However, in general we only require that certain inner products concentrate. We believe this will still hold in the reduced model when one explicitly tracks the form of the error.
7.10 Proof of Corollary 5
@@
8 Miscellaneous Details
8.1 Experimental details
@@
8.2 Semi-real experimental details
8.2.1 Speech data
Speech data was taken directly from Le, (2017) and then was standardized (demeaned and rescaled to have variance 1), as is usual when using ridge-regression.
8.2.2 Residualized returns
@@
8.3 Proof of right-rotational invariance
@@
Matrix normals: Recall that $\boldsymbol{\Sigma}^{\mathrm{col}}$ is drawn from an inverse Wishart distribution with identity scale and $(1+\delta)p$ degrees of freedom; this can be generated by $(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$, where $\mathbf{G}_{2}\in\mathbb{R}^{(1+\delta)p\times p}$ has i.i.d. standard Gaussian entries (since the inverse Wishart has identity scale) (Gelman et al., , 2004, Appendix A). A matrix normal with row covariance $\boldsymbol{\Sigma}^{\text{row}}$ and column covariance $\boldsymbol{\Sigma}^{\text{col}}=(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1}$ can then be generated by $(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}$ (with $\mathbf{G}$ again having i.i.d. Gaussian entries and being independent of $\mathbf{G}_{2}$). The claim then follows from the following. For any fixed $\mathbf{O}\in\mathbb{O}_{p}$,
$(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}(\mathbf{G}_{2}^{\top}\mathbf{G}_{2})^{-1/2}\mathbf{O}$ $\stackrel{{\scriptstyle d}}{{=}}(\boldsymbol{\Sigma}^{\text{row}})^{1/2}\mathbf{G}\mathbf{O}(\mathbf{O}\mathbf{G}_{2}^{\top}\mathbf{G}_{2}\mathbf{O}^{\top})^{-1/2}\mathbf{O}$